factor the following expressions using the method of common factors
8x^2+2xb
x^3y^2+xy

Answer by fractalier(6550)

For 8x^2 + 2xb, both 2 and x are common factors, and we can pull them out, giving us
8x^2 + 2xb = 2x(4x + b)
For (x^3)(y^2) + xy, both x and y are common, and we have
(x^3)(y^2) + xy = xy(x^2y + 1)